OpenStudy (kweenlexis):

-(-2-5x)+(-2)=18 A. x=90 B. x= 18/5 C. - 18/5 D. -90

OpenStudy (moldybubblegum11):

i think its b, but im not sure.

alones (alones):

Hrmmm got B also Sine oyu simplify both sides \(\ −(−2−5x)−2=18\) and simplify this \(\ 5x=18\) than divide both sides by 5 \[\frac{ 5x }{ 5 }=\frac{ 18 }{ 5 }\] which would bring you to \[x= \frac{ 18 }{ 5 }\]

OpenStudy (kweenlexis):

Thank you

OpenStudy (moldybubblegum11):

ya